AnimeGhoul8863:

The function g(n) = n2 − 20n + 95 represents a parabola. Part A: Rewrite the function in vertex form by completing the square. Show your work. (6 points) Part B: Determine the vertex and indicate whether it is a maximum or a minimum on the graph. How do you know? (2 points) Part C: Determine the axis of symmetry for g(n).

AnimeGhoul8863:

@Shadow

AnimeGhoul8863:

Did part A with dude yesterday but now need part B and C

Shadow:

I have to study for a quiz so I won’t be able to help you. Try tagging dude. I know there are easy websites that can help you find the vertex and axis of symmetry. I used them years ago. Also, a minimum just means the vertex is on the bottom, with the lines going up. Maximum means the vertex is the highest point, with the parabola going down.

AnimeGhoul8863:

g(n)= n^2-20n+95 Step 1: n^2 - 20n + 100 Step 2: (b over 2) ^2 So (-20 over 2) ^2=-10^2=100 Step 3: (n^2-20n=100)=95-100 Step 4: (n-10) (n-10)- 5 Step 5: (n-10)^2 - 5 Answer G(n)=(n-10)^2-5

AnimeGhoul8863:

@dude

Shadow:

Do you know what the vertex is?

AnimeGhoul8863:

yes

Shadow:

What is it?

AnimeGhoul8863:

hold on im pulling it up

AnimeGhoul8863:

(10,-5)

Shadow:

Mhm, and have you determined whether this is a minimum or a maximum?

AnimeGhoul8863:

I think its maximum because it goes from negative to positive

Shadow:

What do you mean from negative to positive

AnimeGhoul8863:

hold on imma show u what i have

AnimeGhoul8863:

AnimeGhoul8863:

^^if u look it goes down to up so that would make it maximum because it goes up to the positives

Shadow:

Maximum is when the vertex is the highest point. Does that seem to be the case?

AnimeGhoul8863:

Nu so it would be minimum because the vertex has negatives and it is not at its highest points

Shadow:

Think of it this way: Maximum: Parabola opens down Minimum: Parabola opens up

AnimeGhoul8863:

Ohhhh so Its minimum because the vertext is not at its highest points and becuase the parabola opens up

1 attachment
Shadow:

Yes

AnimeGhoul8863:

*because

Shadow:

And it since it's the lowest point.

AnimeGhoul8863:

how do we know? i know because if the vertex is in the negatives it means its at its lowest points so that would make the parabola open

Shadow:

A vertex can be in the negative and still be a maximum. 

Shadow:

That's why you focus on these two qualifications: Opens up/Opens down Highest point/lowest point It's positioning of the vertex in terms of pos/neg doesn't matter.

AnimeGhoul8863:

so how would you explain how you know if a vertex could still be in negatives but still maximum

Shadow:

Because it's the highest point in the graph and the parabola opens down.

AnimeGhoul8863:

I know because its the lowest point in the graph and the parabola opens up

AnimeGhoul8863:

@dude can yew helps yeet

AnimeGhoul8863:

@dude e.e

AnimeGhoul8863:

Part C im on den one last question

dude:

In order to find axis of symmetry you have to use the this equation: \(\large -\frac{b}{2a}\) Remember: This equation is written as \(g(n)=ax^2+bx+c\)

AnimeGhoul8863:

i know so it would be (n-10) over -5?

AnimeGhoul8863:

cause the equation we finished yesterday is G(n)=(n-10)^2-5

dude:

Ah we're using the original equation for this one \(n^2 − 20n + 95\) As its written in standard form \(ax^2+bx+c\)

AnimeGhoul8863:

im confused now so we dont use our revised equation we use the original

dude:

We use the original

dude:

Its easier to solve for the vertex at least in my opinion

AnimeGhoul8863:

i have the vertex i need the axis symmetry

AnimeGhoul8863:

^^look up i already told shadow the vertex for part B

dude:

OHHHHHHHHH

dude:

Scratch what I just said, just look at the h value Hint: \(g(n)=a(x-h)+k\)

AnimeGhoul8863:

Umm ok.....

dude:

\(g(n)=(n-10)^2-5\) What is the h value?

AnimeGhoul8863:

10?

dude:

Right so x=10 is the axis of symmetry This may help https://www.desmos.com/calculator/s7k0gr4pee

AnimeGhoul8863:

what? u just what IM SOOO CONFUSED how did it go from 10 being h to 10 being x

dude:

The rule for axis of symmetry: \(x=h\) \(g(n)=(x-h)+k\) The value of x for the line of symmetry is equal to the value of h on the vertex form for an equation

AnimeGhoul8863:

so Part C: The Axis equals 10.

dude:

The axis of symmetry is when x=10, yes

AnimeGhoul8863:

Ok are you ready to finish da last question

dude:

give me a few mins, ill be right there. Post it though

AnimeGhoul8863:

ok